### v1

@@ Abstract
Concept drift typically refers to the analysis of changes in data distribution. A drift in the input data can have negative consequences on a learning predictor and the system’s stability. The majority of concept drift methods emphasize the analysis of statistical changes in non-stationary data over time. In this context, we consider another perspective, where the concept drift also integrates substantial changes in the topological characteristics of the data stream. In this article, we introduce a novel framework for monitoring changes in multi-dimensional data streams. We explore a generalization of the standard concept drift focusing on the changes in the topological characteristics of the data. Our developed approach is based on persistent entropy and topology-preserving projections in a continual learning scenario. The framework operates in both unsupervised and supervised environments. To demonstrate the utility of the proposed framework, we analyze the model across three scenarios using data streams generated with MNIST samples. The obtained results reveal the potential of applying topological data analysis for shift detection and encourage further research in this area.
@@ 1 Introduction
In continual learning scenarios, designing a machine learning (ML) model that is robust to distribution shifts is a crucial objective. Traditional ML methods are susceptible to data perturbations, and shifts in input data distribution can significantly affect a model’s performance. Concept drift detectors encompass a family of techniques developed to analyze and detect distribution changes in the context of streaming data and time series. The concept is based on changes in the statistical characteristics of the data over time [1]. In this study, we explore a generalization of the previous definition of concept drift that emphasizes changes in the topological characteristics of the data. There are objects that are essentially equivalent to each other if we consider “equivalence” in the sense that it is possible to define a simple continuous transformation that approximates one object to the other. The essence of an object remains unchanged under simple transformations, such as rotation, translation, scaling, and other types of continuous transformations [2]. On the other hand, there are objects that are essentially different, as it is not feasible to find a continuous transformation to transform one object into another [3]. Or at least, it is not easy to find such a transformation with low computational resources. The field of Topological Data Analysis (TDA), specifically through the mathematical formalism of algebraic topology, defines these concepts of equivalences and differences between objects. Persistent Entropy (PE) is a measure based in Shannon entropy that provides a summary of the geometric information derived from the topological features of a cloud of points [4, 5]. It has been successfully used to effectively distinguish chaotic and periodic time series [5].
@@
In this work, we introduce an extension of the classic concept drift that integrates algebraic topology. We extend the concept of drift, which is based on statistical and geometrical information, with another concept that incorporates the notion of essential sameness and essential difference [3]. Our work aims to provide insight into the research question: Is it adequate to identify a drift between two sequences if there exists a simple continuous bijective function that transforms one sequence into the other? In other words, is it adequate to identify a drift between two sequences when they are equivalent objects in terms of topology? We observe a drift when the significant geometric characteristics of a cloud of points essentially change, becoming different from those of another cloud of points. Figure 1 illustrates examples of objects that can be deformed using simple continuous transformations such as rotation, stretching, bending, and scaling. The first row of the figure displays three digits that can be considered topologically equivalent, while the three objects in the bottom row can also be considered equivalent. However, it is not possible to transform any object in the first row into a digit in the second row. To quantify these types of topological changes, we take into account concepts provided by TDA. We empirically investigate the changes in the distribution of high-dimensional data in an unsupervised scenario, using tools from persistence homology. We develop a general-purpose framework that projects the input data into a low-dimensional space using a dimensionality reduction technique that preserves the topological features of the data. We then apply metrics of persistent homology to evaluate significant changes. The projection is made using Self-Organizing Maps (SOM), which is a mapping technique to reduce dimensionality while preserving the topological characteristics of the input space [6]. In the latent space, we explore the potential of persistent homology to find significant differences among data coming from consecutive chunks. We use the metric of persistent entropy, which summarizes the analysis of persistent homology in a single value [7]. In summary, this work offers the following contributions.
@@ 2 Background
2.1 Concept Drift
Drift detection techniques constitute a family of computational methods for detecting changes in the distribution of time series and data streams. Shifts in data distribution can occur in different forms, probably the most accepted categories are: sudden, gradual, and incremental shifts [8, 9]. A majority of drift detection techniques employ a classifier to categorize incoming instances, and the predictor generates a class label for each input instance, which is then compared to the actual class label. Subsequently, the accuracy is assessed and utilized as a tool to determine whether a drift has occurred. When the classifier’s accuracy significantly decreases, then it is assumed that the data distribution has changed [10, 11, 12]. In this scenario, the effectiveness of various ensemble classifiers has been examined [13, 14, 15, 16]. However, this approach can be applied only in a supervised context, and it requires the presence of ground truth labels, which are not always available. Another set of methods relies on statistical tests directly over the raw data [17, 18, 19, 20]. A lot of distribution shift detection techniques depend on the computation of the empirically estimated distributions. These approaches are sensitive to outliers and noise, and raw data analysis (e.g. applying density estimation) may be affected by the curse of dimensionality [21, 22]. Several studies propose to compare summary of statistics and aggregation metrics of the raw data, for instance, Cumulative Sum and Exponentially Weighted Moving Average [23]. For a more comprehensive review of the latest advancements in the use of data descriptors for concept drift detection, see [24, 25, 26].
2.2 Persistent Entropy
Persistent homology is a key instrument in TDA as it may be used to describe the inherent structure of complex objects such as manifolds [27]. Specifically, persistent homology studies the evolution of $k$-dimensional topological features (often referred to as holes) along a sequence of high-dimensional complex objects (named simplicial complexes[4, 3]. We understand topological features as shapes or data that remain unchanged under certain continuous transformations, such as connected components, independent cycles, and holes [3]. Persistent homology tracks changes of topological features of data that persist across multiple scales, following a specific algorithm that analyzes the connectivity information among the data points [3]. Persistent Entropy, based in Shannon entropy, provides a summary of the information derived from persistent homology [4]. It is a measure for finding significant differences in the geometrical distribution of data points [4, 28]. For a comprehensive and detailed exploration about TDA and persistent homology, see [3, 28].
@@ 3.1 Creation of the latent space
Monitoring and detecting distribution shifts is specially harder in the case of high-dimensional data. Even if some attempts have been introduced in the literature for sparse multivariate time series [29, 30], many algorithms are not properly scalable, for instance the methods based in probability mass distribution. In addition, the norm computation also has limitations in a high-dimensional space (e.g. Euclidean norm) [31]. For this reason, instead of working directly over the points in the original space, we analyze an approach that project the input points into a latent space. The projection is made using dimensionality reduction (DR) techniques, which is a common approach for handling the data in high dimensions. We assume a context that performing a similarity analysis in the original input space may be computationally expensive. Consequently, it is often more resource-efficient to first convert the data into a latent space, and then carry out the similarity analysis. Here, we investigate the projections generated by Self Organizing Maps (SOMs) (also called Kohonen networks) and we compare the results with other two popular DR models (a linear projection (PCA) and the Kernel PCA). The selection of an adequate data descriptor is crucial for ensuring a proper geometry in the latent space preserving the main features of the original space.
SOM is a bio-inspired method that combines concepts from Hebbian learning, vector quantization, and competitive learning [32, 6]. Real-world data most often contain redundancies and inherent correlations among the variables. SOM is a two-layered neural network that transforms intricate relationships among high-dimensional data into straightforward geometric relationships on a standard lattice, typically a two-dimensional grid [33]. Despite its simplicity, the SOM method is effective as a Dimensionality Reduction (DR) technique, a clustering method, and a visualization tool for high-dimensional data [34]. Another advantage is that the method is applicable to unsupervised problems and has the capability to preserve the most important topological features of the reference data [6].
@@ 3.2 Assessing shifts in the latent space
Recently, it was introduced a clustering method based in self-organizing maps for assessing distribution changes in data streams with high dimensional data [35]. Self-organizing map is used for projecting the input data into a latent space, then the analysis is done in the latent space, where the authors computed a distance matrix between the input pattern and cluster centers. The assessment of the distribution shifts is done by applying a statistical summary. This approach of using a data descriptor was also applied in [19, 36], and it is commonly used in methods based on kernel projections [37].
In this work, we modify the analysis based on distances and statistical summaries of the points in the latent space, to an approach that assesses topological changes according to the homological characteristics of the points in the latent space. Once the dimensionality reduction is done, a distance matrix is computed. The distance matrix has the information between the projected point and the cluster centers. Instead of working directly with the coordinates of points in the latent space generated by the DR method, these are summarized using relative locations. Coordinates are arbitrarily selected, and more often than not, they don’t consider any property of the data itself. There are even problems in cases where the coordinates are not natural in any sense [38]. Therefore, a relative location of the latent space points is computed, calculating the distance between the mass centers and the projected points. Hence, our focus is on the geometric properties of the latent space, independent of the chosen coordinates in the latent space. The methodological approach is illustrated in Figure 2. Note that, the approach is general in the sense that can be used any type of DR technique.
@@ 3.3 Pipeline of the proposed framework
  • Dimensionality reduction: When a chunk of data comes, then the input data points are projected using a DR method.
  • Embedding of the geometrical properties in the latent space: For each projected data point, the distance matrix is computed using the data representative of each cluster (centroid distance matrix).
  • Representation of topological features: For each chunk of data, a persistent diagram is computed using the cloud of points created by the centroid distance matrices. Computation of the persistent entropy per chunk (the infinity bar in the computation of persistent entropy is ignored [39].
  • Statistical analysis. A final index is computed between the persistent entropy in one chunk of data and its subsequent chunk. Finally, we generate a sequence of p-value scores computed with the non-parametric Mann-Whitney U test. This sequence of p-value provides the information about significant changes in the topological properties of the data.
Note that an initial training phase is performed to compute the initial clustering and its representative mass centers. An initial time-window of the data stream is used for training the SOM weights and other global parameters. After this initial training phase, we continue the learning process following a usual continual learning scenario.
@@
Figure 3: Generation data streams. Data streams were generated with the MNIST samples interchanged among the different topological types. The graphics illustrate the transition between the sequence of images from one topological type to another type.
@@ 4 Experimental results
In this section, we demonstrate the utility of the proposed approach for monitoring and detecting topological changes in a data stream in the context of continual learning. We designed the experiments to evaluate and contrast the efficacy of the dimensionality reduction methods previously discussed: PCA, Kernel PCA, and SOM. In this ongoing work, we analyze three problems with annotated data streams. There exists a deficiency in the availability of extensive and diverse real-world data streams within high-dimensional spaces for analyzing the impact of distribution changes [8]. This inconvenience is more relevant in the domain of unsupervised analysis of streaming data. As a consequence, we created three synthetic data streams with annotated shifts. The created stream has samples from MNIST [40]. The annotations indicate the time-stamps where was injected the topological differences between two cloud of points.
@@ 4.1 Benchmark data
We generated three synthetic datasets using the MNIST dataset [41], following the methodology outlined in [42, 35, 26]. The procedure consists of creating a stream with chunks of samples that follow a specific distribution, and then alternating these chunks with chunks of instances from a different distribution. By construction, the timing of the distribution changes is predefined. As a consequence, we have the exact time-stamps wherein a drift was injected. We have created the data to check if the method is able to detect “essential” changes between sequences of digits. We divided the digits into three groups: without any holes (without any enclosed space), with one hole, and with two holes. We denote the three experimental studies as: A, B and C. For all the cases, we analyze $20000$ samples. For case studies A and B, the drift is injected every 1000 samples, and for scenario C, the drift occurs every 500 samples.
In case study A, we analyze a data stream where changes occur between chunks with digits in $\{1,3,5,7\}$ and chunks with digits $\{0,6,9\}$ (digits with one hole). Case study B also includes the digits 2 and 4. These digits are problematic due to variations in handwriting. Some people write the numbers 2 and 4 without any hole, while others write them with one hole, it depends on handwriting style. We decided to create the data stream considering the following exchanges between points in $\{1,2,3,5,7\}$ and $\{0,4,6,9\}$. Finally, we also evaluate in the study case C a data stream with the number 8, which is not topologically equivalent to any of the other digits. Case C exchanges samples from the three subsets 1,2,3,5,7, 0,4,6,9, and 8. Figure 3 depicts how the streaming data was generated, in each of the three scenarios. Case study A and B have 20 exchanges between subset of digits with holes and without holes, and case study C has 40 exchanges between subsets of digits without hole, with one hole, and two holes.

### v2

@@ Abstract
In Continual Learning (CL) contexts, concept drift typically refers to the analysis of changes in data distribution. A drift in the input data can have negative consequences on a learning predictor and the system’s stability. The majority of concept drift methods emphasize the analysis of statistical changes in non-stationary data over time. In this context, we consider another perspective, where the concept drift also integrates substantial changes in the topological characteristics of the data stream. In this article, we introduce a novel framework for monitoring changes in multi-dimensional data streams. We explore variations in the topological structures of the data, presenting another angle on the standard concept drift. Our developed approach is based on persistent entropy and topology-preserving projections in a CL scenario. The framework operates in both unsupervised and supervised environments. To show the utility of the proposed framework, we analyze the model across three scenarios using data streams generated with MNIST samples. The obtained results reveal the potential of applying topological data analysis for shift detection and encourage further research in this area.
@@ 1 Introduction
In continual learning scenarios, designing a machine learning (ML) model that is robust to distribution shifts is a crucial objective. Traditional ML methods are susceptible to data perturbations, and shifts in input data distribution can significantly affect the model’s performance. Concept drift detectors encompass a family of techniques developed to analyze and detect distribution changes in the context of streaming data and time series. The concept is based on changes in the statistical characteristics of the data over time [1]. However, there are scenarios where certain modifications in the distributions are not relevant. For example, simple translations or scaling of the data may not provide meaningful information in some contexts. Therefore, it would be beneficial to define concept drift detectors that can detect distribution changes regardless of certain distortions or rotations of the data. There are objects that are essentially equivalent to each other if we consider “equivalence” in the sense that it is possible to define a simple continuous transformation that approximates one object to the other. The essence of an object remains unchanged under simple transformations, such as rotation, translation, scaling, and other types of continuous transformations [2]. On the other hand, there are objects that are essentially different, as it is not feasible to find any continuous transformation to transform one object into another [3]. Or at least, it is not easy to find such a transformation with low computational resources. The field of Topological Data Analysis (TDA), specifically through the mathematical formalism of algebraic topology, defines these concepts of equivalences and differences between objects. Persistent Entropy (PE) is a measure based in Shannon entropy that provides a summary of the geometric information derived from the topological features of a cloud of points [4, 5]. It has been successfully used to effectively distinguish chaotic and periodic time series [5].
@@
In this work, we introduce an extension of the classic concept drift that integrates algebraic topology. We extend the concept of drift, which is based on statistical and geometrical information, with another concept that incorporates the notion of essential sameness and essential difference [3]. Our work aims to provide insight into the research question: Is it adequate to identify a drift between two sequences if there exists a simple continuous bijective function that transforms one sequence into the other? In other words, is it adequate to identify a drift between two sequences when they are equivalent objects in terms of topology? We observe a drift when the geometric characteristics of a point cloud essentially change, becoming different from those of another point cloud. Figure 1 illustrates examples of objects that can be deformed using simple continuous transformations such as rotation, stretching, bending, and scaling. The first row of the figure displays three digits that are topologically equivalent. The second row also shows three topologically equivalent objects. It is not possible to transform any object from the first row into an object of the second row. To quantify these types of topological variations, we take into account concepts provided by TDA. We empirically investigate the changes in the distribution of high-dimensional data in an unsupervised scenario, using tools from persistence homology. We develop a general-purpose framework that projects the input data into a low-dimensional space using a dimensionality reduction technique that preserves the topological features of the data. We then apply metrics of persistent homology to evaluate significant changes. The projection is made using Self-Organizing Maps (SOM), which is a mapping technique to reduce dimensionality while preserving the topological characteristics of the input space [6]. In the latent space, we explore the potential of persistent homology to find significant differences among data coming from consecutive chunks. We use the metric of persistent entropy, which summarizes the analysis of persistent homology in a single value [7]. In summary, this work offers the following contributions.
@@ 2 Background
Drift detection techniques constitute a family of computational methods for detecting changes in the distribution of time series and data streams. Shifts in data distribution can occur in different forms, probably the most accepted categories are: sudden, gradual, and incremental shifts [8, 9]. A majority of drift detection techniques employ a classifier to categorize incoming instances, and the predictor generates a class label for each input instance, which is then compared to the actual class label. Subsequently, the accuracy is assessed and utilized as a tool to determine whether a drift has occurred. When the classifier’s accuracy significantly decreases, then it is assumed that the data distribution has changed [10, 11, 12]. In this scenario, the effectiveness of various ensemble classifiers has been examined [13, 14, 15, 16]. However, this approach can be applied only in a supervised context, and it requires the presence of ground truth labels, which are not always available. Another set of methods relies on empirically estimated distributions and statistical tests directly over the raw data [17, 18, 19, 20]. These approaches may be sensitive to outliers and noise, and raw data analysis (e.g., applying density estimation) may also be affected by the curse of dimensionality [21, 22]. Several studies propose to compare summary of statistics and aggregation metrics of the raw data, for instance, Cumulative Sum and Exponentially Weighted Moving Average [23]. For a more comprehensive review of the latest advancements in the use of data descriptors for concept drift detection, see [24, 25, 26].
Persistent homology is a key instrument in TDA as it may be used to describe the inherent structure of complex objects such as manifolds [27]. Specifically, persistent homology studies the evolution of $k$-dimensional topological features (often referred to as holes) along a sequence of high-dimensional complex objects (named simplicial complexes[4, 3]. We understand topological features as shapes or data that remain unchanged under certain continuous transformations, such as connected components, independent cycles, and holes [3]. This process can be seen as tracking changes across filtrations at multiple scales, following a specific algorithm that analyzes the connectivity information among the data points [3]. Persistent Entropy, based in Shannon entropy, provides a summary of the information derived from persistent homology [4]. It is a measure for finding significant differences in the geometrical distribution of data points [4, 28]. For a comprehensive and detailed exploration about TDA and persistent homology, see [3, 28].
@@ 3.1 Creation of the latent space
Monitoring and detecting distribution shifts is specially harder in the case of high-dimensional data. Although some attempts have been introduced in the literature for sparse multivariate time series [29, 30], the scalability of existing algorithms remains an issue. In particular, methods based on probability mass distribution encounter significant challenges in high-dimensional spaces. In addition, the computation of distances between vectors also has limitations in a high-dimensional space (e.g. Euclidean norm) [31], and a similarity analysis in the original input space may be computationally expensive. Consequently, it is often more resource-efficient to first convert the data into a latent space, and then carry out the similarity analysis. For this reason, a common approach is to transform the input data into a latent space, instead of making the analysis directly in the original space. We analyze an approach that projects the input points into a latent space using dimensionality reduction (DR) techniques, which is a common method for handling data in high dimensions. Here, we investigate the projections generated by Self Organizing Maps (SOMs) (also called Kohonen networks), and we compare the results with other two popular DR models (a linear projection (PCA) and the Kernel PCA). The selection of an adequate data descriptor is crucial for ensuring a proper geometry in the latent space preserving the main features of the original space.
SOM is a bio-inspired method that combines concepts from Hebbian learning, vector quantization, and competitive learning [32, 6]. Real-world data most often contain redundancies and inherent correlations among the variables. SOM is a two-layered neural network that transforms intricate relationships among high-dimensional data into straightforward geometric relationships on a standard lattice, typically a two-dimensional grid [33]. Despite its simplicity, the SOM model is effective as a DR method, a clustering method, and a visualization tool for high-dimensional data [34]. Another advantage is that the method is applicable to unsupervised problems and has the capability to preserve the most important topological features of the reference data [6].
@@ 3.2 Assessing shifts in the latent space
Recently, it was introduced a clustering method based in SOM for assessing distribution changes in data streams with high dimensional data [35]. SOM is used for projecting the input data into a latent space, then the analysis is done in the latent space, where the authors computed a distance matrix between the input pattern and cluster centers. The assessment of the distribution shifts is done by applying a statistical summary. This approach of using a data descriptor was also applied in [19, 36], and is commonly used in methods based on kernel projections [37]. Here, we modify the framework introduced in  [35], which is based on distances and statistical summaries of the points in the latent space, to an approach that assesses topological changes according to the homological characteristics of the points in the latent space. Once the DR mapping is done, a distance matrix is computed. The distance matrix has the information between the projected point and the cluster centers. We use relative locations instead of directly working with the coordinates of points in the latent space generated by the DR method. The coordinates are arbitrarily selected and often do not consider any property of the data itself. There are even problems in cases where the coordinates are not natural in any sense [38]. Therefore, the relative locations of the point cloud in the latent space are computed by calculating the distances between the mass centers and the projected points. Hence, our focus is on the geometric properties of the latent space, independent of the chosen coordinates in the latent space. This methodological approach is illustrated in Figure 2. Note that the approach is general in the sense that any type of DR technique can be used.
@@ 3.3 Pipeline of the proposed framework
  • i)
    Initial setup: An initial batch of points is used as a pre-phase procedure to train the selected DR method.
  • ii)
    Dimensionality reduction: When a chunk of data comes, then the input data points are projected using the chosen DR method.
  • iii)
    Embedding of the geometrical properties in the latent space: For each projected data point, the distance between the projected point and each centroid is computed. By centroid, we refer to the mass center of a cluster, which can be thought of as the average position of all the points within the cluster, or a representative data point that best exemplifies the cluster’s characteristics. Then, for each chunk, a matrix is created (with dimensions of number of clusters $\times$ Chunk size) where the column vector has the distances between a projected point and the centroids.
  • iv)
    Representation of topological features: A persistent diagram is created using the distance matrix described in the previous step. The persistence diagram summarizes the information in the distance matrix (note that each projected point is computed a distance matrix). Then, the persistent entropy is computed from the collection of persistence diagrams within the chunk, ignoring the infinity bar [39]. The representation of topological features returns a real vector with a number of elements equal to the size of the chunk.
  • v)
    Statistical analysis. We compute a final index that corresponds to a comparison between the representation of topological features in the current chunk and the representation of topological features in a reference chunk. In our experiments, we compare two consecutive chunks. The comparison between these two vectors is done using a non-parametric test. In our experiments, we applied the Mann-Whitney U test. However, analysis of similarity between these two vectors can be done using other tools. The statistical test provides a p-value score, which we use for monitoring changes between a current chunk of data and a reference chunk of data.
Finally, when the global procedure is applied online, it generates a sequence of p-value scores. This sequence provides information about significant changes in the topological properties of the data stream. Note, an initial training phase is performed to compute the initial clustering and its representative mass centers. An initial time window of the data stream is used for training the SOM weights and other global parameters. After this initial training phase, we continue the learning process following a usual continual learning scenario. The procedure described previously can include an additional step that involves adjusting the DR technique once a shift is detected. When a shift is detected, the SOM method can be re-trained using an arbitrary batch of data (for example, the last chunk or the last $n$ chunks of data), with the precautions to avoid catastrophic forgetting. This online calibration procedure was not investigated in this initial empirical exploration and could be a promising direction for future research.
@@
Figure 3: Creation of synthetic case studies. Data streams were generated with the MNIST samples interchanged among the different topological types. The graphics illustrate the transition between the sequence of images from one topological type to another type.
@@ 4 Experimental results
In this section, we explore the utility of the proposed approach for monitoring and detecting topological changes in a data stream in the context of continual learning. We designed the experiments to evaluate and contrast the efficacy of the dimensionality reduction methods previously discussed: PCA, Kernel PCA, and SOM. There exists a deficiency in the availability of extensive and diverse real-world data streams within high-dimensional spaces for analyzing the impact of distribution changes [8]. This inconvenience is more relevant in the domain of unsupervised analysis of streaming data. As a consequence, we created three synthetic data streams with annotated shifts. Our analysis is unsupervised (we do not use the data labels); we use the information about the annotated shifts to evaluate the performance of the proposed approach. The created stream has samples from MNIST [40]. The annotations indicate the time stamps where the topological differences between two cloud of points were injected.
@@ 4.1 Benchmark data
We generated three synthetic datasets using the MNIST dataset [41], following the methodology outlined in [42, 35, 26]. The procedure consists of creating a stream with chunks of samples that follow a specific distribution, and then alternating these chunks with chunks of instances from a different distribution. By construction, the time stamps of the distribution changes are arbitrary predefined. As a consequence, we have marked the exact time stamps at which a drift was injected to evaluate the capacity of the approach in monitoring shifts. We created the data to check if the method can detect “essential” changes between sequences of digits. We divided the digits into three groups: those without any holes (zero-dimensional homology), those with one hole (one-dimensional homology), and those with two holes (one-dimensional homology with two loops). We denote the three experimental studies as: A, B and C. For the three cases, we analyze $20000$ samples. For case studies A and B, the drift is injected every 1000 samples, and for scenario C, the drift occurs every 500 samples. In case study A, we analyze a data stream where the changes occur between chunks with digits in $\{1,3,5,7\}$ (zero-dimensional homology) and chunks with digits in $\{0,6,9\}$ (one-dimensional homology). Case study B also includes the digits 2 and 4. These digits are problematic due to variations in handwriting. Some individuals write the numbers 2 and 4 without any hole, while others write them with one hole, depending on individual handwriting style. Then,case study B has a data stream considering the following exchanges between points in $\{1,2,3,5,7\}$ and $\{0,4,6,9\}$. Finally, in case study C, we evaluate a data stream that includes the number 8, which is not topologically equivalent to any of the other digits. Case C exchanges samples from the three subsets $\{1,2,3,5,7\}$, $\{0,4,6,9\}$, and $\{8\}$. Figure 3 depicts how the streaming data was generated for each of the three scenarios. Case study A and B have 20 exchanges between subsets containing digits with and without holes, and case study C involves 40 exchanges between subsets containing digits without holes, with one hole, and two holes.